How many grams of dry NH4Cl need to be added to 2.50 L of a 0.700 M solution of ammonia, NH3, to prepare a buffer solution that has a pH of 8.62? Kbfor ammonia is 1.8×10−5.

Express your answer with the appropriate units.
Answer #1

pH = pKb + log10([salt]/[base])

[base] = 2.5 * 0.7 = 1.75 M

pkb = -log10(1.8*10-5) = 5.7447

14 - 8.62 = 4.7447 + log([salt]/1.75)

log([salt]/1.75) = 0.6353

[salt] = 7.5568 M

[salt] = mass/volume = mass/2.5 = 7.5568

mass of salt to be added = 18.892 g.

Hope this helps!